What is the molarity of a solution with a volume of 5 L made with 600g of NaOH?


1
Expert's answer
2021-03-31T08:02:20-0400

Molarity= Mole of the Solute/Volume of Solution

Mole of Solute=Mass/Molar Mass of NaOH

Molar Mass of NaOH=40g/mole

Mole=600g÷40g/mole= 15Mol

Molarity=15÷5= 3M

Molarity = 3M